# Robust $\mathcal{H}_{\infty}$ control and worst-case search in constrained parametric space

### v5

@@ Abstract
Standard $\mathcal{H}_{\infty}$$\mathcal{H}_{2}$ robust control and analysis tools operate on uncertain parameters assumed to vary independently within prescribed bounds. This paper extends their capabilities in the presence of constraints coupling these parameters and restricting the parametric space. Focusing on the worst-case search, we demonstrate – based on the theory of upper-$C^{1}$ functions – the validity of standard, readily available smooth optimization to address this nonsmooth constrained optimization problem. Specifically, we prove that the sequential quadratic programming algorithm converges to Karush-Kuhn-Tucker points, and that such conditions are satisfied by any subgradient at a local minimum. This worst-case search then enables robust controller synthesis: identified worst-case configurations are iteratively added to an active set on which a non-smooth multi-models optimization of the controller is performed. From a practical point of view, we combine the local exploitation proposed above with a global exploration using Monte-Carlo sampling. The proposed approach is illustrated through the robust control of a mechanical system, of order 50 with 43 uncertain parameters. We show that the constrained optimization effectively complements Monte-Carlo sampling by enabling fast detection of rare worst-case configurations, and that the robust controller optimization converges with less than 10 active configurations.
I Introduction
The $\mathcal{H}_{\infty}$ norm is a widely used metric to quantify the performance of control systems. For a linear system, it is equivalently defined as the $\mathcal{L}_{2}$-induced norm, hence representing the maximum energy amplification from input to output, or as the maximum peak in the frequency domain:
@@ I Introduction
where $||\cdot||$ may refer to the $\mathcal{H}_{2}$ or $\mathcal{H}_{\infty}$ system norm, and the transfer function $T_{zw}(\mathrm{s},K,\delta)$ is now written as a function of $K$ the controller to be optimized, $\mathcal{K}$ is the set of controllers of the chosen structure, and $\delta$ is the vector of $k$ uncertain parameters in a parametric space $\mathcal{D}$. In presence of multiple control requirements, the following formulation may be used:
@@
Classically in the literature, $\mathcal{D}$ is the hypercube $1,1]^{k}$ after normalization of the uncertain parameters. In this context, an overview of this control problem and of the modern resolution method is presented in [10]. The goal of this paper is to address the presence of nonlinear constraints restricting this hypercube.
We mainly focus on the inner maximization problems – it will then allow us to solve Problems (1) and (2) by leveraging state-of-art nonsmooth optimization of the controller [8]. Namely, search for worst-case performance:
@@
and worst-case stability – stability is implicitly required to ensure finite norms – treated as in [3]:
@@
where $A(K,\delta)$ is the matrix A of the state-space representation of the closed loop, and $\alpha(\cdot)$ denotes the spectral abscissa:
When the parametric space is the whole hypercube $1,1]^{k}$, Matlab’s function systune, relying on the non-smooth optimization of [8, 3], is now widely used in academia and industry to solve problems (1) and (2). More specifically, in [3], problems (3) and (4) are solved with local non-smooth optimization. Indeed, the computation of Clarke subdifferentials [21] of the spectral abscissa and of the $\mathcal{H}_{\infty}$ norm [16] enables a bundle method to quickly find worst-case configurations with regard to stability and performance. The control problem (1) is then tackled by combining this worst-case search with the non-smooth controller optimization of [8]. This approach relies on the construction of an inner approximation $\mathcal{D}_{a}\subset\mathcal{D}$ composed of so-called active configurations. At each iteration, the controller is tuned so as to be robust to all current active configurations. Then, a worst-case search is performed; if the worst-case configuration does not satisfy the control requirements, it is added to $\mathcal{D}_{a}$ and the algorithm goes back to controller tuning; otherwise, the algorithm can terminate. It is shown that this approach is both more reliable and more efficient than gridding the parametric space, which quickly becomes intractable when the number of parameters increases. In practice, the optimization generally finishes with a small number of active configurations. As this approach relies on local searches, robustness is generally thoroughly verified a posteriori. $\mu$-analysis [27] was developed for this purpose, generally based on branch-and-bound algorithms [36]. It provides guaranteed deterministic lower and upper bounds of the global optimum; a probabilistic variant also exists [13, 38]. Monte-Carlo sampling can also guarantee a probability of robustness within predefined thresholds of accuracy and confidence [40]. Another common industrial practice is to grid the parametric space with a limited number of samples. But this approach has exponential complexity with respect to the number of parameters, will miss any worst case occurring between grid points and does not provide any statistical guarantee.
@@
The robust control and analysis approaches discussed above do not address the constrained case, at least not in their current implementations. Our focus is twofold. Firstly, we seek computationally efficient techniques that can be integrated into robust control design, alternating between multi-models controller synthesis and worst-case search as in [3], to solve the standard problem (1) and its multi-objectives formulation (2) in reasonable time. Secondly, we aim to assess their reliability in consistently detecting global worst-case configurations as a validation tool.
@@
Specifically, we demonstrate new properties of the class of upper-$C^{1}$ functions [39], and in particular the convergence of Sequential Quadratic Programming (SQP) [29], a standard constrained optimization algorithm, when applied to such functions even though they are not differentiable everywhere. We deduce that SQP can be applied to the nonsmooth problem (3), and to (4) under an additional assumption, to find their Karush-Kuhn-Tucker points. From a practical point of view, we also study the use of Monte-Carlo sampling to perform a global exploration of the parameter space, with SQP subsequently ensuring efficient local exploitation. Then, following the approach of [3], this worst-case search method is combined with multi-models controller optimization [8] to perform robust control. The proposed approach is illustrated through the control of a mechanical system, which typically exhibits constraints of the form (5). To the best of our knowledge, this is the first robust $\mathcal{H}_{\infty}$ control approach that tackles nonlinear constraints on the uncertain parameters.
@@
We emphasize that the proposed approach, although studied here in the context of the constrained space (5), may also prove valuable in the unconstrained case. Indeed, $\mu$-analysis remains difficult to apply to industrial systems with many uncertain parameters. By contrast, the proposed approach scales well to plants with many states and parameters (as a gradient-based method), albeit without formal guarantee of global optimality – this problem being known to be NP-hard. Furthermore, the nonsmooth optimization of [3], implemented in systune, is local and sometimes misses worst-case configurations, whereas the proposed approach performs a global exploration of the parametric space.
II Constrained optimization of upper-$C^{1}$ functions
II-A Properties of upper-$C^{1}$ functions
In this section, we introduce the class of upper-$C^{1}$ functions and demonstrate favorable properties for their minimization. This will be relevant for the robust control problem, which, as we shall see, involves such functions. For readability and since our primary focus is on robust control, proofs that result from adapting existing arguments from the differentiable to the upper-$C^{1}$ functions are only provided in Appendix.
Definition 1 (Lower- and upper-$C^{1}$ functions [39])
@@ II-A Properties of upper-$C^{1}$ functions
[3, Sections III.A., II.B., V.A.] shows that minimization of such upper-$C^{1}$ functions, even when not differentiable everywhere, exhibits similar behavior to smooth functions. In particular, the authors show the convergence of their bundle method, in the search of the worst-case $\mathcal{H}_{\infty}$ norm, even when the cutting plane step is reduced to a standard linesearch (as in smooth optimization). This work follows [25] that treats of the optimization of lower- and upper-$C^{1}$ functions. Similar considerations are also thoroughly discussed by the same authors in [11]. Similarly, in [26], the Frank-Wolfe algorithm, originally designed for smooth functions, is also proven to converge to Clarke-stationary points when applied to upper-$C^{1}$ functions. In the same spirit, we now show that the upper-$C^{1}$ functions, although not differentiable everywhere, maintain some favorable properties for their minimization. Then, we will demonstrate the convergence of one particular smooth optimization algorithm when applied to upper-$C^{1}$ functions.
In this whole section, we consider $f:\mathbb{R}^{k}\rightarrow\mathbb{R}$ an upper-$C^{1}$ function. Let us first consider the unconstrained problem:
$\underset{x\in\mathbb{R}^{k}}{\text{minimize}}\;f(x)$ (6)
We recall that the directional derivative at $x$ in direction $d$ of $f$ defined as
$f^{\prime}(x,d)=\lim_{t\to 0^{+}}\frac{f(x+td)-f(x)}{t}$
always exists for an upper-$C^{1}$ function [35]. The subdifferential of $f$ at $x$ is denoted $\partial f(x)$.
Proposition 1 (Descent directions and optimality of upper-$C^{1}$ functions – Unconstrained case)
@@
b) Optimality. If $x$ locally minimizes $f(x)$, then $g_{x}||=0$ for any $g_{x}\in\partial f(x)$.
Proof.
a) [26, Proposition 2 and equation (6)] shows that the directional derivative of an upper-$C^{1}$ function $f$ is:
@@
Proposition 1.a) shows that when a smooth optimization algorithm is provided with any subgradient as if it were the gradient of a smooth function, progress can still be achieved at points of nonsmoothness, owing to the regularity property of upper-$C^{1}$ functions. Naturally, the selected direction is not necessarily the steepest descent. Proposition 1.b) ensures that an optimization algorithm can rely on the same optimality conditions as in the smooth case when picking any subgradient (we recall that, for the general case of nonsmooth functions, the optimality condition is $0\in\partial f(x)$ which is a weaker condition, also called Clarke-stationarity, and requires a specific treatment in nonsmooth optimization algorithms).
@@
$\begin{array}[]{lll}&\underset{x\in\mathbb{R}^{k}}{\text{minimize}}&f(x)\\ &\text{subject to }&c_{i}(x)\leq 0\;,\;\;1\leq i\leq n\end{array}$ (8)
@@
We also consider the quadratic problem:
$Q(x,H,g_{x}):\;\begin{array}[]{lll}&\underset{p\in\mathbb{R}^{k}}{\text{minimize}}&g_{x}^{T}p+\frac{1}{2}p^{T}Hp\\ &\text{subject to }&c_{i}(x)+\nabla c_{i}(x)^{T}p\leq 0\;,\;\;1\leq i\leq n\end{array}$ (10)
@@
Proposition 2 (Descent directions and optimality of upper-$C^{1}$ functions – Constrained case)
@@
Proof.
a) Cf. Appendix. The proof is a mere adaptation of [28, Theorem 3.1] to the upper-$C^{1}$ case.
@@
$g_{x}=-\sum_{j=1}^{p}a_{s_{j}}\sum_{i=1}^{n}\lambda_{i}(s_{j})\nabla c_{i}(x)=-\sum_{i=1}^{n}\left(\sum_{j=1}^{p}a_{s_{j}}\lambda_{i}(s_{j})\right)\nabla c_{i}(x)=-\sum_{i=1}^{n}\mu_{i}\nabla c_{i}(x)$
where $\mu_{i}:=\sum_{j=1}^{p}a_{s_{j}}\lambda_{i}(s_{j})$. For all $i$: $\mu_{i}\geq 0$ (because all $a_{s_{j}}$ and $\lambda_{i}(s_{j})$ are $\geq 0$), and $\mu_{i}c_{i}(x)=0$ (because if $c_{i}(x)\neq 0$, then all $\lambda_{i}(s_{j})$ are 0). This proves Proposition 2.b). ∎
@@
Proposition 1.b) is also proved in [26, Proposition 4]. Proposition 2.a) is a generalization of [28, Theorem 3.1], originally for differentiable functions, to upper-$C^{1}$ functions. To the best of our knowledge, Proposition 2.b) is novel, and can also be seen as a specific case of the usual KKT conditions for upper-$C^{1}$ functions. It allows optimality to be checked with any subgradient, in contrast to the general nonsmooth case, where only the existence of one subgradient satisfying KKT conditions is required.
II-B Convergence of sequential quadratic programming (SQP) for upper-$C^{1}$ functions
We now introduce the Sequential Quadratic Programming (SQP) algorithm to solve Problem (8). It is described in Algorithm 1, as in the original reference [28] for differentiable functions, except that we replaced the gradient $\nabla f(x)$ by any subgradient $g_{x}\in\partial f(x)$ in the quadratic problem $Q(x,H,g_{x})$ (Eq. (10)). See also [33] for a discussion when the penalty parameter $r$ is variable. We do not detail how the quadratic problem in Step 1 or the step size $\lambda_{k}$ in Step 2 may be solved, or how $H_{k}$ is updated in Step 3, as it does not affect the convergence proof (as long as the conditions described in Algorithm 1 are satisfied)
Algorithm 1 Sequential Quadratic Programming (SQP) algorithm [28] for upper-$C^{1}$ functions
Initialization: $x_{0}\in\mathbb{R}^{k}$, $H_{0}\in\mathbb{R}^{k\times k}$ positive definite, $r>0$, $\lambda_{max}>0$.
Step 1: Pick a $g_{x_{k}}\in\partial f(x_{k})$. Find a KKT point $p_{k}$ of the quadratic problem $Q(x_{k},H_{k},g_{x_{k}})$ defined in Eq. (10).
Step 2: Set $x_{k+1}=x_{k}+\lambda_{k}p_{k}$ for any $\lambda_{k}\in[0,\lambda_{max}]$ satisfying:

### v6

# Worst-case search in constrained uncertainty space for robust $\mathcal{H}_{\infty}$ synthesis

@@ Abstract
Standard $\mathcal{H}_{\infty}$$\mathcal{H}_{2}$ robust control and analysis tools operate on uncertain parameters assumed to vary independently within prescribed bounds. This paper extends their capabilities in the presence of constraints coupling these parameters and restricting the parametric space. Focusing on the worst-case search, we demonstrate – based on the theory of upper-$C^{1}$ functions – the validity of standard, readily available smooth optimization to address this nonsmooth constrained optimization problem. Specifically, we prove that for such functions, any subgradient satisfy Karush-Kuhn-Tucker (KKT) conditions at a local minimum, and that any accumulation point of the sequential quadratic programming (SQP) is a KKT point. From a practical point of view, we combine this local exploitation with a global exploration using Monte-Carlo sampling. This worst-case search then enables robust controller synthesis: identified worst-case configurations are iteratively added to an active set on which a non-smooth multi-models optimization of the controller is performed. The proposed approach is illustrated through the robust control of a mechanical system. We show that this method enables fast detection of rare worst-case configurations, and that the robust controller optimization converges with a limited number of active configurations.
keywords:
Robust control, uncertain systems, nonsmooth optimization
thanks: Corresponding author.
1 Introduction
The $\mathcal{H}_{\infty}$ norm is a widely used metric to quantify the performance of control systems. For a linear system, it is equivalently defined as the induced 2-norm, hence representing the maximum energy amplification from input to output, or as the maximum peak in the frequency domain:
@@ 1 Introduction
where $||\cdot||$ may refer to the $\mathcal{H}_{2}$ or $\mathcal{H}_{\infty}$ system norm, and the transfer function $T_{zw}(\mathrm{s},K,\delta)$ is now written as a function of $K$ the controller to be optimized, $\mathcal{K}$ is the set of controllers of the chosen structure, and $\delta$ is the vector of $k$ uncertain parameters in a parametric space $\mathcal{D}$. Multiple control requirements are generally addressed with the following formulation:
@@
Classically in the literature, $\mathcal{D}$ is the hypercube $1,1]^{k}$ after normalization of the uncertain parameters. In this context, an overview of this control problem and of the modern resolution method, relying on nonsmooth optimization, is presented in [9]. The goal of this paper is to address the presence of nonlinear constraints restricting this hypercube.
We mainly focus on the inner maximization problem, often called worst-case search; solving it will then allow us to treat Problems (1) and (2) by leveraging state-of-art nonsmooth optimization of the controller [8]. Namely, the worst-case search in performance:
@@
and the worst-case search in stability, implicitly required to ensure finiteness of the norms, treating it as in [3]:
@@
where $A(K,\delta)$ is the matrix A of the state-space representation of the closed loop, and $\alpha(\cdot)$ denotes the spectral abscissa.
When the parametric space is the whole hypercube $1,1]^{k}$, Matlab’s function systune, relying on the non-smooth optimization of [8, 3], is now widely used in academia and industry to solve problems (1) and (2). More specifically, in [3], problems (3) and (4) are solved with local non-smooth optimization. Indeed, the computation of Clarke subdifferentials [20] of the spectral abscissa and of the $\mathcal{H}_{\infty}$ norm [15] enables a bundle method to quickly find worst-case configurations with regard to stability and performance. The control problem (1) is then tackled by combining this worst-case search with the non-smooth controller optimization of [8]. This approach relies on the construction of an inner approximation $\mathcal{D}_{a}\subset\mathcal{D}$ composed of so-called active configurations. At each iteration, the controller is tuned so as to be robust to all current active configurations. Then, a worst-case search is performed; if the worst-case configuration does not satisfy the control requirements, it is added to $\mathcal{D}_{a}$ and the algorithm goes back to controller tuning; otherwise, the algorithm can terminate. It is shown that this approach is both more reliable and more efficient than gridding the parametric space, which quickly becomes intractable when the number of parameters increases. In practice, the optimization generally finishes with a small number of active configurations. As this approach relies on local searches, robustness is generally thoroughly verified a posteriori. The $\mu$-analysis [26] was developed for this purpose, generally based on branch-and-bound algorithms [34]. It provides guaranteed deterministic lower and upper bounds of the global optimum; a probabilistic variant also exists [12, 36].
@@
The robust control and analysis approaches discussed above do not address the constrained case, at least not in their current implementations. Our focus is twofold. Firstly, we seek computationally efficient techniques that can be integrated into robust control design, alternating between multi-models controller synthesis and worst-case search as in [3], to solve Problems (1) and (2) in reasonable time. Secondly, we aim to assess their reliability in consistently detecting global worst-case configurations as a validation tool.
Problem (3) is widely seen as a non trivial task, due to the nonsmoothness of the $\mathcal{H}_{\infty}$ norm. Nevertheless, [3] links the $\mathcal{H}_{\infty}$ norm, as a function of a structured controller or uncertainty, to the theory of lower- and upper-$C^{1}$ functions [37], a class of functions with a variety of regularity properties [33]. Multiple references suggest that the minimization of upper-$C^{1}$ functions, even when not differentiable everywhere, exhibits similar behavior to smooth functions. In [3], the authors show the convergence of their bundle method even when the cutting plane step is reduced to a standard linesearch (as in smooth optimization) when applied to such functions. This work follows [24] that treats of the optimization of lower- and upper-$C^{1}$ functions. Similar considerations are also thoroughly discussed by the same authors in [10] in the context of robust control. In [25], the Frank-Wolfe algorithm, originally designed for smooth functions, is also proven to converge to Clarke-stationary points when applied to upper-$C^{1}$ functions (in the context of unconstrained optimization).
@@
In this paper, we demonstrate new properties of the class of upper-$C^{1}$ functions, and in particular that any subgradient verifies Karush-Kuhn-Tucker (KKT) conditions when the constraint functions $c_{i}$ are differentiable. We also propose a slight adaptation of the Sequential Quadratic Programming (SQP) [28], a smooth constrained optimization algorithm, to the upper-$C^{1}$ case, and demonstrate the convergence to KKT points. We deduce that this algorithm can be applied to the nonsmooth problem (3), and to (4) under an additional assumption. From a practical point of view, we also take advantage of Monte-Carlo sampling and its stochastic properties [38] to perform a global exploration of the parameter space (in the stochastic sense), with SQP subsequently ensuring efficient local exploitation. Then, following the approach of [3], this worst-case search method is combined with multi-models controller optimization [8] to perform robust control, with the possibility to add multiple worst-case configurations to the active set at each iteration. The proposed approach is illustrated through the control of a realistic mechanical system, which typically exhibits constraints of the form (5).
@@
We emphasize that the proposed approach, although studied here in the context of the constrained space (5), may also prove valuable in the unconstrained case. Indeed, $\mu$-analysis remains difficult to apply to industrial systems with many uncertain parameters, since the exact computation of the structured singular value is NP-hard [39], forcing reliance on bounds that are often either intractable to compute or overly conservative. By contrast, the proposed approach scales well to plants with many states and parameters (as a gradient-based method), albeit without formal guarantee of global optimality. Furthermore, the nonsmooth optimization of [3], implemented in systune, is local and sometimes misses worst-case configurations, whereas the proposed approach performs a global exploration in the stochastic sense; it also opens the possibility to speed up the optimization by adding multiple active configurations at each iteration.
Notations
Given $T_{zw}(s)$ a transfer function between an input $w$ and an output $z$, $T_{zw}(s)||$ may denote either the $\mathcal{H}_{\infty}$ or the $\mathcal{H}_{2}$ norm for the sake of generality; however, the paper will mostly focus on the $\mathcal{H}_{\infty}$ norm (because the $\mathcal{H}_{2}$ norm is differentiable, and thus does not cause any particular difficulty). $\alpha(\cdot)$ denotes the spectral abscissa:
Given $f:\mathbb{R}^{n}\to\mathbb{R}$ a locally Lipschitz function, the Clarke subdifferential of $f$ at point $x$ is denoted $\partial f(x)$. When defined, the directional derivative of $f$ at $x$ in direction $d$ is defined as
$f^{\prime}(x,d)=\lim_{t\to 0^{+}}\frac{f(x+td)-f(x)}{t}\;.$
2 Constrained optimization of upper-$C^{1}$ functions
2.1 Properties of upper-$C^{1}$ functions
In this section, we introduce the class of upper-$C^{1}$ functions and demonstrate favorable properties for their minimization. As we shall see in Section 3 in the context of robust $\mathcal{H}_{\infty}$ control, the worst-case search in performance (3), and in stability (4) under an additional assumption, can be addressed with such functions.
Definition 1.
: Lower- and upper-$C^{1}$ functions [37]
@@ 2.1 Properties of upper-$C^{1}$ functions
In this whole section, we consider $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ an upper-$C^{1}$ function. We recall that the directional derivative always exists for an upper-$C^{1}$ function [33]. Let us first consider the unconstrained problem:
$\underset{x\in\mathbb{R}^{n}}{\text{minimize}}\;f(x)$ (6)
Proposition 2.
: Descent directions and optimality of upper-$C^{1}$ functions – Unconstrained case
@@
b) Optimality. If $x$ locally minimizes $f(x)$, then $g_{x}||=0$ for any subgradient $g_{x}\in\partial f(x)$.
{pf}
a) [25, Proposition 2 and equation (6)] shows that the directional derivative of an upper-$C^{1}$ function $f$ is:
@@
Proposition 2.a) shows that when a smooth optimization algorithm is provided with any subgradient as if it were the gradient of a smooth function, progress can still be achieved at points of nonsmoothness. Naturally, the selected direction is not necessarily the steepest descent. Proposition 2.b) ensures that an optimization algorithm can rely on the same optimality conditions as in the smooth case when picking any subgradient (we recall that, for the general case of nonsmooth functions, the optimality condition is $0\in\partial f(x)$ which is a weaker condition, also called Clarke-stationarity, and requires a specific treatment in nonsmooth optimization algorithms).
@@
$\begin{array}[]{lll}&\underset{x\in\mathbb{R}^{n}}{\text{minimize}}&f(x)\\ &\text{subject to }&c_{i}(x)\leq 0\;,\;\;1\leq i\leq n\end{array}$ (8)
@@
We also define the quadratic problem $Q(x,H,g_{x})$:
$\begin{array}[]{lll}&\underset{p\in\mathbb{R}^{n}}{\text{minimize}}&g_{x}^{T}p+\frac{1}{2}p^{T}Hp\\ &\text{subject to }&c_{i}(x)+\nabla c_{i}(x)^{T}p\leq 0\;,\;\;1\leq i\leq n\end{array}$ (10)
@@
Proposition 3.
Descent directions and optimality of upper-$C^{1}$ functions – Constrained case
@@
{pf}
a) Cf. Appendix. The proof is a mere adaptation of [27, Theorem 3.1] to the upper-$C^{1}$ case.
@@
$g_{x}$ $=-\sum_{j=1}^{p}a_{s_{j}}\sum_{i=1}^{n}\lambda_{i}(s_{j})\nabla c_{i}(x)$
$=-\sum_{i=1}^{n}\left(\sum_{j=1}^{p}a_{s_{j}}\lambda_{i}(s_{j})\right)\nabla c_{i}(x)$
$=-\sum_{i=1}^{n}\mu_{i}\nabla c_{i}(x)$
where $\mu_{i}:=\sum_{j=1}^{p}a_{s_{j}}\lambda_{i}(s_{j})$. For all $i$: $\mu_{i}\geq 0$ (because all $a_{s_{j}}$ and $\lambda_{i}(s_{j})$ are $\geq 0$), and $\mu_{i}c_{i}(x)=0$ (because if $c_{i}(x)\neq 0$, then all $\lambda_{i}(s_{j})$ are 0). This proves Proposition 3.b).
@@
Proposition 2.b) is also proved in [25, Proposition 4]. Proposition 3.a) is a generalization of [27, Theorem 3.1], originally for differentiable functions, to upper-$C^{1}$ functions. It will be used in next section to prove convergence of the optimization algorithm. To the best of our knowledge, Proposition 3.b) is novel, and can also be seen as a specific case of the usual KKT conditions for upper-$C^{1}$ functions. It allows optimality to be checked with any subgradient, in contrast to the general nonsmooth case, where only the existence of one subgradient satisfying KKT conditions is required.
2.2 Sequential quadratic programming (SQP) algorithm for upper-$C^{1}$ functions
We now introduce the Sequential Quadratic Programming (SQP) algorithm adapted for upper-$C^{1}$ functions to solve Problem (8). It is described in Algorithm 1, and is identical to that of the original reference [27] for differentiable functions, except that we replaced the gradient $\nabla f(x)$ by any subgradient $g_{x}\in\partial f(x)$ in the quadratic problem $Q(x,H,g_{x})$ (Eq. (10)). We do not detail how the quadratic problem in Step 1 or the step size $\lambda_{k}$ in Step 2 may be solved, or how $H_{k}$ is updated in Step 3, as it does not affect the convergence proof (as long as the conditions described in Algorithm 1 are satisfied).
Algorithm 1 Sequential Quadratic Programming (SQP) algorithm [27] adapted for upper-$C^{1}$ functions
Initialization: $x_{0}\in\mathbb{R}^{n}$, $H_{0}\in\mathbb{R}^{n\times n}$ positive definite, $r>0$, $\lambda_{max}>0$.
Step 1: Pick any $g_{x_{k}}\in\partial f(x_{k})$. Find a KKT point $p_{k}$ of the quadratic problem $Q(x_{k},H_{k},g_{x_{k}})$.
Step 2: Set:
$x_{k+1}=x_{k}+\lambda_{k}p_{k}$
and find a $\lambda_{k}\in[0,\lambda_{max}]$ satisfying: